# jina-embeddings-v4: Universal Embeddings for Multimodal Multilingual Retrieval

### v2

@@ 7 Evaluation
jina-embeddings-v4 (dense) 72.19 84.11 84.11 66.49 55.97 71.59 67.11 72.70 85.89
jina-embeddings-v4 (late) 79.29 90.17
@@
jina-clip-v2 39.84 53.61 81.12
colpali-v1.2 (late) 64.50 83.90
dse-qwen2-2b-mrl-v1 (dense) 65.96 85.80
@@ A.1 Datasets in the Jina-VDR Benchmark
jinaai/arabic_chartqa_ar Mixed Charts Question 745 / 745 ar
@@
jinaai/ChartQA Mixed Charts Question 7996 / 1000 en
@@
jinaai/docqa_energy Energy Digital Docs Question 69 / 972 en
@@
jinaai/docqa_healthcare_industry Medial Digital Docs Question 90 / 963 en
@@
jinaai/europeana-nl-legal Legal Scans Question 199 / 300 nl
jinaai/github-readme-retrieval-multilingual† Software / IT Markdown Docs Description 16755 / 4398 ar, bn, de, en, es, fr, hi, id, it, ja, ko, nl pt, ru, th, vi, zh
jinaai/hindi-gov-vqa Governmental Digital Docs Question 454 / 340 hi
jinaai/hungarian_doc_qa_hu Mixed Digital Docs Question 54 / 54 hu
@@
jinaai/owid_charts_en Mixed Charts Question 132 / 972 en
@@
jinaai/wikimedia-commons-documents-ml† Mixed Mixed Description 14061 / 14661 ar, bn, de, en, es, fr, hi, hu, id, it, ja, ko, my, nl, pt, ru, th, ur, vi, zh
@@
jinaai/stanford_slide Education Slides Question 14 / 1000 en
@@
jinaai/table-vqa Science Tables Question 992 / 1000 en
jinaai/tatqa Finance Digital Docs Question 121 / 176 en
jinaai/tqa Education Illustrations Question 981 / 394 en
@@
jinaai/wikimedia-commons-maps Mixed Maps Description 443 / 455 en
@@ A.2 JinaVDR (Visual Document Retrieval) Benchmark Results
Average 45.48 48.14 40.21 65.51 48.62 73.86 80.16
medical-prescriptions 38.18 37.25 15.66 83.91 38.22 81.17 97.95
stanford_slide 81.78 95.28 91.48 100 100 100 100
donut_vqa 19.39 2.60 1.63 32.53 25.31 78.48 73.55
table-vqa 35.64 34.24 21.06 54.66 57.39 58.90 60.91
ChartQA 26.01 29.46 36.45 50.29 55.01 66.45 67.40
tqa 50.04 24.21 27.45 68.27 65.29 65.79 68.28
openai-news 76.63 87.30 70.05 94.82 92.56 93.75 97.62
europeana-de-news 11.26 12.05 11.19 34.64 44.75 49.05 65.65
europeana-es-news 51.99 44.03 13.14 44.74 60.58 60.10 80.58
europeana-it-scans 39.11 38.69 16.23 54.32 53.92 57.88 73.14
europeana-nl-legal 34.97 29.07 9.79 30.89 29.50 37.14 54.15
hindi-gov-vqa 1.83 7.52 5.02 13.04 9.80 15.40 21.94
automobile_catalogue_jp 20.92 49.53 32.54 39.69 65.62 77.10 81.32
beverages_catalogue_ru 11.05 14.11 39.45 60.36 81.07 85.31 86.65
ramen_benchmark_jp 28.03 62.19 39.79 47.46 51.75 89.50 94.65
jdocqa_jp_ocr 1.64 7.79 19.91 39.45 66.73 75.57 82.34
hungarian_doc_qa 32.43 55.42 47.91 73.08 53.80 71.19 75.20
arabic_chartqa_ar 6.27 5.91 4.12 12.03 35.37 44.73 49.79
arabic_infographicsvqa_ar 13.26 13.49 49.98 51.24 71.93 84.86 94.07
owid_charts_en 65.50 61.23 57.61 83.53 83.48 87.40 89.68
arxivqa 56.67 54.27 84.64 95.44 93.33 95.44 95.44
docvqa 76.37 43.87 45.23 88.84 86.28 84.05 92.98
shiftproject 59.43 68.43 32.10 76.56 78.76 81.65 90.72
docqa_artificial_intelligence 90.78 82.93 64.81 95.19 97.52 96.43 98.04
docqa_energy 87.52 78.68 65.39 94.95 90.08 89.19 96.57
docqa_gov_report 86.29 82.03 69.18 94.75 94.50 91.84 95.97
docqa_healthcare_industry 84.83 88.92 67.77 95.85 95.24 94.62 97.51
tabfquad 42.62 78.03 46.94 87.62 91.80 95.57 94.79
mpmqa_small 85.54 66.85 59.53 88.95 81.99 80.96 91.80
jina_2024_yearly_book 87.67 85.98 76.63 93.05 92.90 94.78 97.68
wikimedia-commons-maps 5.37 5.12 20.43 30.53 33.14 39.55 54.15
plotqa 61.13 51.39 23.98 73.46 75.98 77.97 78.77
MMTab 74.82 74.18 44.63 85.80 85.90 86.16 90.00
CharXiv-en 46.85 41.58 56.34 83.96 84.10 82.80 87.78
student-enrollment 1.05 1.19 0.72 3.92 4.08 7.88 12.02
tatqa 75.21 70.38 44.91 85.29 80.68 79.46 92.93
shanghai_master_plan 12.69 93.32 75.28 90.50 93.21 94.04 97.18
europeana-fr-news 24.55 23.78 16.21 30.99 37.49 36.10 50.44
infovqa 68.64 75.39 63.03 90.21 92.41 92.11 96.49
@@
Average 21.95 37.34 48.32 46.82 58.43 65.79 74.50
Arabic (ar) 19.60 38.06 45.29 41.96 62.92 72.07 81.19
Bengali (bn) 22.93 44.44 48.48 38.56 52.39 67.31 76.60
German (de) 12.74 39.45 52.72 51.91 63.00 70.32 80.69
English (en) 36.44 45.21 56.51 68.76 70.31 73.91 81.72
Spanish (es) 12.75 46.02 54.70 59.57 66.31 71.80 80.33
French (fr) 15.59 35.83 35.69 44.42 40.84 53.67 59.28
Hindi (hi) 16.53 36.69 47.47 38.13 50.56 63.45 68.78
Hungarian (hu) 25.32 33.86 44.20 26.22 52.60 65.36 76.10
Indonesian (id) 28.78 39.35 50.77 56.41 61.59 66.34 73.59
Italian (it) 19.58 37.93 49.61 49.08 60.12 64.18 73.31
Japanese (jp) 21.32 30.42 43.75 53.50 63.61 66.72 76.63
Korean (ko) 34.86 35.16 47.12 54.77 68.47 71.83 81.53
Burmese (my) 22.84 29.57 54.25 16.31 36.97 46.91 50.23
Dutch (nl) 14.89 39.76 50.23 56.87 64.67 68.97 78.42
Portuguese (pt) 23.27 45.57 53.97 58.17 67.26 69.23 78.43
Russian (ru) 16.82 38.92 49.34 46.77 64.26 69.56 80.45
Thai (th) 30.00 29.79 45.76 47.90 55.78 63.36 70.03
Urdu (ur) 13.64 32.78 36.30 15.68 38.49 48.71 61.91
Vietnamese (vi) 32.40 39.87 54.06 56.82 64.33 72.93 80.18
Chinese (zh) 18.76 28.18 46.24 54.62 64.14 69.11 80.55
@@
Average 50.11 65.17 38.93 76.83 72.26 85.52 85.69
Arabic (ar) 27.49 27.99 31.07 58.53 55.63 75.04 75.71
Bengali (bn) 1.29 28.29 27.06 53.72 47.55 65.60 66.03
German (de) 60.11 84.50 43.35 87.16 80.76 91.10 91.39
English (en) 87.40 91.60 48.58 93.91 90.75 96.78 97.35
Spanish (es) 78.57 83.30 43.19 86.84 78.61 89.46 89.99
French (fr) 77.55 83.66 42.12 85.42 79.20 90.29 90.31
Hindi (hi) 2.72 48.19 28.53 57.93 46.62 68.98 70.88
Indonesian (id) 78.05 82.41 38.59 83.06 74.39 88.31 88.69
Italian (it) 78.83 86.53 44.33 86.94 80.93 91.74 91.36
Japanese (jp) 14.46 63.12 41.99 75.99 75.31 89.60 90.91
Korean (ko) 40.01 35.32 38.32 69.45 68.77 86.92 86.85
Dutch (nl) 76.52 86.33 43.17 88.01 82.97 92.83 91.39
Portuguese (pt) 80.33 84.54 43.86 86.83 80.22 91.58 91.50
Russian (ru) 39.78 51.07 36.80 80.67 79.08 89.55 88.63
Thai (th) 1.47 36.75 37.80 68.68 65.51 77.78 76.46
Vietnamese (vi) 66.70 79.65 37.08 75.00 68.15 86.93 87.13
Chinese (zh) 40.52 54.57 35.94 67.98 73.90 81.43 82.08
@@
Average 22.29 42.69 55.36 78.32 62.81 64.05 85.46
Arabic (ar) 0.38 1.67 49.26 77.02 52.51 52.86 78.11
German (de) 48.27 66.85 52.29 75.70 57.50 64.22 85.97
English (en) 51.38 63.61 48.28 78.70 63.31 63.87 85.17
Spanish (es) 54.28 63.80 53.16 77.81 62.38 63.61 84.24
French (fr) 51.69 64.53 54.92 75.95 62.48 62.20 85.35
Hindi (hi) 0.08 0.08 89.06 96.51 96.62 87.14 96.86
Hungarian (hu) 15.55 61.71 52.05 73.50 58.36 63.75 85.70
Japanese (jp) 0.40 47.64 54.76 70.38 58.03 61.25 85.61
Russian (ru) 0.47 2.83 47.09 78.10 57.51 63.18 82.95
Chinese (zh) 0.45 54.16 52.78 79.51 59.07 58.44 84.66
@@
Average 7.20 1.12 2.12 16.45 11.06 8.13 37.42
Arabic (ar) 1.10 0.39 0.47 2.59 3.26 2.28 6.63
German (de) 4.03 0.71 5.85 30.58 14.46 9.32 42.26
English (en) 48.39 1.72 4.73 47.51 13.07 13.06 63.48
Spanish (es) 6.25 0.18 1.93 24.15 8.46 9.12 40
French (fr) 3.86 1.92 2.03 18.02 12.21 8.55 29.69
Hindi (hi) 0.16 0.78 0.84 1.81 4.51 3.96 17.56
Hungarian (hu) 5.58 0.63 3.10 23.09 11.74 6.87 27.21
Japanese (jp) 0.36 1.70 0.46 4.12 14.71 7.65 46.21
Russian (ru) 1.67 1.39 0.79 11.18 13.98 8.43 40.43
Chinese (zh) 0.58 1.77 1.04 1.49 14.17 12.09 60.70

### v3

@@ 7 Evaluation
jina-embeddings-v4 (dense) 73.98 84.11 84.11 66.49 55.97 71.59 67.11 72.70 85.89
jina-embeddings-v4 (late) 80.55 90.17
@@
jina-clip-v2 40.52 53.61 81.12
colpali-v1.2 (late) 63.80 83.90
dse-qwen2-2b-mrl-v1 (dense) 67.25 85.80
@@ A.1 Datasets in the Jina-VDR Benchmark
jinaai/arabic_chartqa_ar Mixed Charts Question 745 / 342 ar
@@
jinaai/ChartQA Mixed Charts Question 996 / 834 en
@@
jinaai/docqa_energy Energy Digital Docs Question 69 / 971 en
@@
jinaai/docqa_healthcare_industry Medial Digital Docs Question 90 / 961 en
@@
jinaai/europeana-nl-legal Legal Scans Question 199 / 244 nl
jinaai/github-readme-retrieval-multilingual† Software / IT Markdown Docs Description 16953 / 16998 ar, bn, de, en, es, fr, hi, id, it, ja, ko, nl pt, ru, th, vi, zh
jinaai/hindi-gov-vqa Governmental Digital Docs Question 454 / 337 hi
jinaai/hungarian_doc_qa_hu Mixed Digital Docs Question 54 / 51 hu
@@
jinaai/owid_charts_en Mixed Charts Question 132 / 937 en
@@
jinaai/wikimedia-commons-documents-ml† Mixed Mixed Description 15593 / 15217 ar, bn, de, en, es, fr, hi, hu, id, it, ja, ko, my, nl, pt, ru, th, ur, vi, zh
@@
jinaai/stanford_slide Education Slides Question 14 / 994 en
@@
jinaai/table-vqa Science Tables Question 992 / 385 en
jinaai/tatqa Finance Digital Docs Question 121 / 270 en
jinaai/tqa Education Illustrations Question 981 / 393 en
@@
jinaai/wikimedia-commons-maps Mixed Maps Description 443 / 451 en
@@ A.2 JinaVDR (Visual Document Retrieval) Benchmark Results
Average 46.88 48.97 40.96 65.39 68.89 75.47 81.52
medical-prescriptions 38.18 38.12 15.68 66.22 38.86 80.95 97.69
stanford_slide 81.78 95.28 91.48 100.0 100.0 100.0 97.16
donut_vqa 19.39   2.59   1.46 34.12 25.31 78.60 74.08
table-vqa 55.22 63.04 36.34 80.98 85.70 86.57 89.21
ChartQA 28.39 31.47 39.73 54.45 58.38 70.88 71.80
tqa 50.11 24.40 27.80 63.03 65.35 65.44 68.46
openai-news 76.63 87.30 70.05 94.81 93.75 93.97 96.43
europeana-de-news 11.26 12.02 11.18 35.20 44.32 48.89 63.76
europeana-es-news 51.99 43.82 12.95 45.70 60.66 60.81 80.70
europeana-it-scans 39.11 38.77 16.54 58.70 54.28 58.01 73.29
europeana-nl-legal 39.38 34.24 11.30 39.13 33.12 42.77 59.82
hindi-gov-vqa   1.83   7.51   5.21 11.43 10.19 15.32 22.49
automobile_catalogue_jp 20.92 50.39 32.54 35.72 66.44 72.22 81.32
beverages_catalogue_ru 11.05 14.09 39.66 68.47 80.32 85.68 87.73
ramen_benchmark_jp 28.02 63.37 41.28 52.03 51.66 90.77 94.65
jdocqa   1.64   7.85 19.94 35.68 67.00 75.63 82.42
hungarian_doc_qa 34.28 57.84 50.44 68.83 55.25 74.64 75.56
arabic_chartqa_ar   9.32   8.63   6.62 26.92 49.35 62.16 66.64
arabic_infographicsvqa_ar 13.26 13.43 50.36 34.76 71.72 85.38 93.21
owid_charts_en 66.19 62.10 57.71 78.17 84.26 92.06 92.29
arxivqa 56.73 54.41 83.41 92.54 93.33 95.44 95.44
docvqa 81.11 50.81 45.29 90.38 86.28 83.06 92.98
shiftproject 62.42 70.25 31.85 75.18 78.54 82.55 91.13
docqa_artificial_intelligence 91.68 82.98 66.52 96.09 97.52 96.43 98.04
docqa_energy 89.97 76.97 65.56 96.03 90.08 88.66 96.28
docqa_gov_report 87.20 82.72 68.84 92.92 94.19 92.03 95.97
docqa_healthcare_industry 86.44 86.88 68.13 93.14 96.14 94.62 97.51
tabfquad 45.67 80.49 47.04 89.18 92.38 95.57 95.38
mpmqa_small 85.54 67.39 59.72 88.88 81.62 80.44 91.28
jina_2024_yearly_book 87.67 85.98 77.12 95.77 93.39 94.29 98.17
wikimedia-commons-maps   5.37   5.06 20.67 27.46 33.06 40.23 53.45
plotqa 61.13 51.44 24.05 70.58 75.99 77.48 78.75
MMTab 74.82 74.06 44.54 84.66 86.04 86.08 90.03
CharXiv-en 46.85 41.47 56.28 79.64 83.86 83.00 87.66
student-enrollment   1.05   1.30   0.70   3.95   4.09   8.04 11.55
tatqa 75.62 49.88 44.23 82.57 80.97 80.14 92.76
shanghai_master_plan 12.69 92.67 75.28 88.87 92.56 95.53 97.41
europeana-fr-news 24.55 23.69 16.43 30.33 38.23 36.66 50.16
infovqa 73.61 75.09 63.38 87.53 92.64 92.16 96.69
@@
Average 21.99 37.43 48.63 33.60 58.67 66.04 75.63
Arabic (ar) 19.62 38.40 45.85 28.40 63.06 71.41 81.81
Bengali (bn) 22.93 44.55 49.37 26.63 52.89 66.98 76.41
German (de) 12.74 39.58 52.87 40.36 62.99 70.21 80.86
English (en) 36.45 45.24 56.58 64.98 70.23 73.55 81.66
Spanish (es) 12.75 46.10 54.85 41.34 66.43 71.68 80.82
French (fr) 15.59 36.06 35.73 43.93 41.32 53.58 59.42
Hindi (hi) 16.73 36.94 48.42 18.02 50.94 62.64 71.77
Hungarian (hu) 25.38 33.88 44.42 12.67 52.35 65.86 76.00
Indonesian (id) 28.79 39.48 50.85 40.46 62.03 66.02 73.72
Italian (it) 19.63 37.98 49.77 34.76 60.05 63.96 73.68
Japanese (jp) 21.41 30.43 44.03 28.83 63.71 66.50 77.13
Korean (ko) 34.98 35.24 47.61 29.82 68.37 71.45 81.77
Burmese (my) 22.84 29.45 54.36 10.28 37.61 56.58 65.01
Dutch (nl) 14.90 39.89 50.40 52.29 65.09 68.58 78.94
Portuguese (pt) 23.32 45.85 54.28 51.30 67.53 69.04 78.85
Russian (ru) 16.82 38.95 49.34 31.88 64.44 68.86 80.70
Thai (th) 30.00 29.64 46.25 39.13 56.41 61.68 71.02
Urdu (ur) 13.64 32.73 36.52   9.45 38.76 49.76 62.17
Vietnamese (vi) 32.40 39.80 54.59 43.72 64.62 73.30 80.24
Chinese (zh) 18.82 28.41 46.45 23.82 64.51 69.23 80.58
@@
Average 50.11 65.14 39.06 72.91 72.24 85.57 85.69
Arabic (ar) 27.49 27.98 31.02 55.19 55.95 75.02 75.26
Bengali (bn)   1.29 28.27 26.96 49.25 47.30 65.70 66.08
German (de) 60.11 84.58 45.46 84.15 80.62 91.09 91.35
English (en) 87.43 91.67 48.69 91.10 90.69 96.94 97.34
Spanish (es) 78.57 83.31 43.35 84.02 78.70 89.60 90.19
French (fr) 77.55 83.54 42.42 83.73 79.11 90.25 90.45
Hindi (hi)   2.72 48.08 28.55 51.22 46.49 69.31 70.98
Indonesian (id) 78.05 82.46 38.59 79.67 74.57 88.42 88.62
Italian (it) 78.83 86.54 44.26 85.31 80.81 91.76 91.41
Japanese (jp) 14.46 63.20 42.02 69.02 75.42 89.74 90.80
Korean (ko) 40.01 35.23 37.87 64.16 68.83 87.04 86.89
Dutch (nl) 76.52 86.36 43.25 84.10 82.85 92.83 91.37
Portuguese (pt) 80.33 84.46 43.88 85.00 80.09 91.43 91.47
Russian (ru) 39.78 50.86 37.04 78.16 78.92 89.51 88.61
Thai (th)   1.47 36.67 37.62 65.29 65.45 77.61 76.67
Vietnamese (vi) 66.70 79.67 37.14 70.05 68.20 86.90 86.94
Chinese (zh) 40.52 54.53 35.89 60.05 74.05 81.44 82.26
@@
Average 22.30 42.77 55.36 76.36 62.76 78.10 85.34
Arabic (ar)   0.38   1.67 49.36 77.31 52.73 66.15 77.66
German (de) 48.27 66.86 52.49 73.53 57.35 79.38 85.63
English (en) 51.38 63.66 48.35 77.13 63.47 77.92 85.36
Spanish (es) 54.28 63.44 53.44 79.02 62.57 78.68 84.62
French (fr) 51.69 64.76 54.94 76.91 62.17 78.65 85.27
Hindi (hi)   0.08   0.08 88.55 93.39 97.00 97.46 96.50
Hungarian (hu) 15.55 62.31 52.30 71.06 58.17 80.09 85.01
Japanese (jp)   0.40 47.80 54.74 70.00 57.76 77.04 85.67
Russian (ru)   0.47   3.07 47.08 70.72 57.43 76.33 83.11
Chinese (zh)   0.45 54.04 52.30 74.54 58.94 69.33 84.55
@@
Average   7.20   1.13   2.13 10.42 11.10   8.18 37.51
Arabic (ar)   1.10   0.40   0.47   3.06   3.64   2.20   6.20
German (de)   4.03   0.71   5.54 20.17 15.09   9.27 41.94
English (en) 48.39   1.70   4.83 23.26 12.94 13.33 64.17
Spanish (es)   6.25   0.18   2.10 18.06   8.61   9.11 39.84
French (fr)   3.86   2.00   2.05 10.86 11.87   8.70 30.55
Hindi (hi)   0.16   0.86   0.82   3.19   4.93   4.05 17.44
Hungarian (hu)   5.58   0.69   3.01   7.34 11.10   6.69 27.30
Japanese (jp)   0.36   1.53   0.54   3.44 14.91   7.63 45.65
Russian (ru)   1.67   1.39   0.88 13.16 13.61   8.66 40.80
Chinese (zh)   0.58   1.84   1.04   1.62 14.28 12.14 61.19